On July 1, 2016, Killearn Company acquired 105,000 of the outstanding shares of Shaun Company for $19 per share. This acquisition gave Killearn a 40 percent ownership of Shaun and allowed Killearn to significantly influence the investee's decisions. As of July 1, 2016, the investee had assets with a book value of $5 million and liabilities of $1,101,500. At the time, Shaun held equipment appraised at $280,000 above book value; it was considered to have a seven-year remaining life with no salvage value. Shaun also held a copyright with a five-year remaining life on its books that was undervalued by $650,000. Any remaining excess cost was attributable to goodwill. Depreciation and amortization are computed using the straight-line method. Killearn applies the equity method for its investment in Shaun. Shaun's policy is to declare and pay a $1 per share cash dividend every April 1 and October 1. Shaun's income, earned evenly throughout each year, was $640,000 in 2016, $670,600 in 2017, and $735,200 in 2018. In addition, Killearn sold inventory costing $121,800 to Shaun for $203,000 during 2017. Shaun resold $104,500 of this inventory during 2017 and the remaining $98,500 during 2018.

1. Determine the equity income to be recognized by Killearn during each of these years.

2. Compute Killearn's investment in Shaun Company's balance as of December 31, 2018.
Answer #1

Questiion: 1. Determine the equity income to be recognized by Killearn during each of these years.

Ans. Equity Income for the 17 will be 40% of $670600=$268240-Dividend Recd. (@ 1+1=2*105000)210000= $ 58240

Dividend received for the year will be deducted from the Investment cost as per Equity Method used.

Equity Income for the 18 will be 40% of $735200=$294080-Dividend Recd. (@ 1+1=2*105000)210000= $ 84080

Dividend received for the year will be deducted from the Investment cost as per Equity Method used.

NOTE: The Inventory transactions profits will be taken direct effect in Profit & Loss Statement as usual.

Question 2. Compute Killearn's investment in Shaun Company's balance as of December 31, 2018.

Ans. Original Investment at Cost is $ 105000*19=$1995000

Add profit shared (Net of Dividend) 17= $58240(2017)

Add profit shares (Net of Dividend) 18= $84080(2018)

Investment Value shown in 2018 would be (30th June 2018)=$2137320

Note: The assumptions taken in answer above is that July to June is accounting years.

None of the effect has been taken for part of the year ie from 1st july 18 to 31/12/2018. otherwise we may deduct dividend for $1 (half year) from Investment value. ie $2137320-105000=2032320, but as the profits are not avaialble for part of the year, the inestment cost how much reduced with our profit sharing cant be possible to calculated.,